OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 1,024 and a4 = -16? I'm not sure how to find the r without knowing the second term

OpenStudy (anonymous):

( 1024 + (-16) )/4, u get 260

OpenStudy (anonymous):

i mean -260 which is ur r

OpenStudy (anonymous):

ur 6th term would then be -16-260-260 which will be -536

OpenStudy (anonymous):

so then would i take -16 and minus 260 from it 2 more times?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

hint: a2 = a1*r a2 = 1024r ------------ a3 = a2*r a3 = 1024r*r a3 = 1024r^2 ------------ a4 = a3*r a4 = 1024r^2*r a4 = 1024r^3 -16 = 1024r^3 now solve for r

OpenStudy (anonymous):

uhm, that actually is not one of the answers x3 @ 11 calc Select one: a. 1 b. -0.25 c. -1 d. 0.25 and @ jim, I'm not understanding.

jimthompson5910 (jim_thompson5910):

basically each term is found by multiplying the previous term by the common ratio r

jimthompson5910 (jim_thompson5910):

the second term a2 is found by multiplying the previous term a1 by r that's why a2 = a1*r

jimthompson5910 (jim_thompson5910):

and why a3 = a2*r

OpenStudy (anonymous):

Yeah, I understand that. But the only thing i could think to do is to divide the a1 by 2, but I can tell that's not what i need to do

jimthompson5910 (jim_thompson5910):

a1 does NOT mean "a times 1" a2 does NOT mean "a times 2" etc so you can't pull out that 2 and divide by it

jimthompson5910 (jim_thompson5910):

do you see how I got -16 = 1024r^3 ?

OpenStudy (anonymous):

uhm yes i think i understand, so a2 would be 1024r? so a16 would be 1024r^15?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

so because a4 = 1024r^3, and a4 = -16, we know that a4 = 1024r^3 -16 = 1024r^3 -16/1024 = r^3 -1/64 = r^3 r^3 = -1/64 r = cube root of (-1/64) r = -1/4 once you know this value of r, you can use it to find any term you want

OpenStudy (anonymous):

Okay, thank you! I got -1, is that correct?

jimthompson5910 (jim_thompson5910):

let me check

OpenStudy (anonymous):

Oh i already answered it and it was right ^-^ thank you for your help

jimthompson5910 (jim_thompson5910):

yes -1 is correct, and I'm glad it agrees, you're welcome